OpenStudy (anonymous):

write a differential equation for the statement: the rate of change of y with respect to x is proportional to the square of the sum of x and 6. Medal and fan for sure:)

OpenStudy (anonymous):

dy/dx = k (x+6)^2, where k is the proportional constant

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=k(x+6)^2\]

OpenStudy (anonymous):

oh wow i was overthinking it thanks a bunch guys :)

OpenStudy (anonymous):

its a proportion so theres a constant k in there

OpenStudy (anonymous):

Yea that's right. Thanks bro